# Same Task, Different Circuits: Disentangling Modality-Specific Mechanisms in VLMs

### v2

@@ 1 Introduction
Recent years have seen major progress in building vision-and-language models (VLMs) that can handle both image and text inputs (Li et al., 2025). However, a significant performance gap persists between textual tasks and their visual counterparts (Fu et al., 2024; Wang et al., 2024a; Cohen et al., 2024). For example, models are considerably less accurate at counting objects in an image than at counting words in a list (Figure 1a, Appendix C), or at identifying a winner in a board game when it is supplied as an image rather than as a textual description (Fu et al., 2024). Curiously, this performance gap persists even in VLMs with a decoder that is jointly trained from scratch on both modalities (e.g. Team et al. (2024)). To address the performance gap between analogous image and text tasks in VLMs, one should first understand its source.
@@
In this paper, we aim to explain the performance gap in VLMs between analogous visual and textual tasks from structural and functional perspectives. First, we create a dataset of five diverse tasks, each comprising pairs of aligned textual and visual prompt variants. Each prompt consists of data, query, and generation positions, which contain the subject of the prompt (an image or a short text), the task description, and the last token position, respectively. For example, the data for an object counting task can be a text listing objects (“car flower banana…”) or an image with these objects, and the query for this task will ask “How many bananas are there?”. We use this dataset to extract and examine the circuits—the computational sub-graphs—responsible for performing each of the variants. We measure the degree of structural overlap between these modality-specific circuits, and find that vision and language tasks are performed by relatively disjoint VLM components (Figure 1a), with an average of only 18% components shared between modalities.
@@
To analyze the functional differences of intermediate sub-circuits, we swap them between the two modalities, and measure how interchangeable they are for task performance (Figure 1b). This reveals two main findings: first, despite being structurally disjoint, the sub-circuit that processes the query tokens, as well as the sub-circuit that generates the answer, are functionally equivalent between modalities—leading to similar performance when swapped. Second, there is a major functional difference in the processing of data tokens between modalities, resulting in much lower performance after swapping the relevant sub-circuit. Hence, we conclude that the difference in data processing is the leading cause behind the accuracy gap.
@@
We utilize these findings to bridge the performance gap, without any additional training, by automatically intervening in the model’s computation at test time. Specifically, we patch the visual token embeddings from later layers of the model—where these embeddings are more aligned to their textual counterparts—back to earlier layers (Figure 1c). This results in visual processing that is more aligned with analogous textual prompts, leading to higher accuracy on visual prompts. Our results suggest that understanding the differences in the mechanisms between modalities is attainable, and provides a path to improving visual processing in VLMs.
@@ 2 Preliminaries
In this section we present our main unit of analysis—model circuits. We discuss how circuits for specific tasks are discovered and evaluated. These discovery and evaluation techniques are used in Section 4 to identify the common and distinct sub-circuits for analogous tasks in different modalities.
@@ 2.1 Model Circuits
A circuit (Elhage et al., 2021) is a minimal subset of interconnected model components that perform the computations required for a specific task or prompt. In our analysis, we consider a circuit component to be either an entire attention head or an individual multilayer perceptron (MLP) neuron, at a specific output position. We follow the existing definition of MLP neurons (Geva et al., 2021; Nikankin et al., 2024), where the $n^{\text{th}}$ neuron in an MLP block is the combination of the $n^{\text{th}}$ row vector of the first MLP layer and the $n^{\text{th}}$ column vector of the second MLP layer. A full definition for MLP neurons is given in Appendix A.
@@ 2.2 Circuit Discovery
To identify circuits for both textual and visual task variants, we employ causal analysis techniques that score the importance of each component’s activation (Vig et al., 2020). For each textual or visual task, we divide its prompts into two subsets: one for discovery and another for evaluation. From the discovery subset, for each prompt $p$ with answer $r$, we select a random counterfactual prompt $p^{\prime}$ that yields a different answer $r^{\prime}$. Ideally, we would find the importance of a component $u$ for a prompt $p$ by intervening on its activation $a_{u,p}$ with the counterfactual activation $a_{u,p^{\prime}}$, and calculate $\operatorname{LD}(r,r^{\prime}|\operatorname{do}(a_{u,p}=a_{u,p^{\prime}}))$, the difference in logits of $r$ and $r^{\prime}$ post-intervention (Wang et al., 2022). However, this process becomes unfeasible due to the large amount of model components. Instead, we approximate this effect by applying attribution patching (Nanda, 2022) with integrated gradients (AP-IG; Hanna et al., 2024):
$\operatorname{E}(u,p,p^{\prime})=\left|\left(a_{u,p^{\prime}}-a_{u,p}\right)\frac{1}{k}\sum_{i=1}^{k}\frac{\partial\operatorname{LD}\left(r,r^{\prime}|\operatorname{do}\left(e=e^{\prime}+\frac{i}{k}\left(e-e^{\prime}\right)\right)\right)}{\partial a_{u,p}}\right|$ (1)
where $k=5$ (Hanna et al., 2024) and $e$ and $e^{\prime}$ are the input embeddings for the prompts $p$ and $p^{\prime}$, respectively. This value is a first-order Taylor-series approximation to the logit difference, and can be calculated once, for all model components in parallel. The value is averaged across prompts and calculated per circuit component. A higher value indicates greater component importance for the analyzed task prompts. Given a score for each component, we construct a circuit $\mathbf{c}$ by selecting a specific percentage of the highest-scoring components.
@@ 2.3 Circuit Evaluation
We evaluate $\mathbf{c}$ by measuring its faithfulness (Wang et al., 2022)—the proportion of the full model’s task performance explained by the circuit. For each prompt $p$ with answer $r$ in the evaluation subset, we randomly choose a counterfactual prompt $p^{\prime}$ with answer $r^{\prime}$ and compute the activations of the model for $p^{\prime}$. We then pass $p$ through the model while ablating all non-circuit components using the pre-computed counterfactual activations, as per standard procedure (Mueller et al., 2025). The intervention’s effect is quantified by the logit difference between correct and counterfactual answers, denoted $\operatorname{LD}(r,r^{\prime}|\operatorname{do}(\forall u\notin\mathbf{c}:a_{u,p}=a_{u,p^{\prime}}))$. This effect is normalized to calculate the faithfulness of circuit $\mathbf{c}$ on task $\mathrm{T}$ (Mueller et al., 2025):
$\operatorname{F}(\mathbf{c},\mathrm{T})=\frac{1}{|\mathrm{T}|}\sum_{(p,p^{\prime},r,r^{\prime})\in\mathrm{T}}\frac{\operatorname{LD}(r,r^{\prime}|\operatorname{do}(\forall u\notin\mathbf{c}:a_{u,p}=a_{u,p^{\prime}}))-\operatorname{LD}_{\mathbf{M}}(r,r^{\prime})}{\operatorname{LD(r,r^{\prime})}-\operatorname{LD}_{\mathbf{M}}(r,r^{\prime})}$ (2)
@@ 3 Experimental Settings
To investigate the accuracy gap between textual and visual prompts in VLMs, we construct a dataset of five question-answering tasks, seen in Figure 2. Each task consists of a query paired with data presented in one of two analogous formats: as either an image or a text. This analogous design ensures direct comparability—for every image-based prompt, we create an analogous text-based prompt that describes the same information. Our dataset contains the following tasks:
@@
To perform circuit discovery on these tasks, we enforce two constraints on the prompts. First, we only use prompts where the answer is a single word. Second, for consistency within a task, we align all prompts to a positional template, such that each position in each prompt always contains the same type of token in it. (For instance, in factual recall, the first four tokens contain the entity name.) Prompts that do not align well to the template are filtered out. These constraints are standard in circuit discovery (Vig et al., 2020; Mueller et al., 2025). Prompt examples per task and further details on prompt generation are provided in Section B.2.
@@
To ensure we identify meaningful circuits within these models, we follow standard procedures (Mueller et al., 2025) and verify high VLM performance (substantially above-chance accuracy) on each task. We report accuracies in Appendix C and focus on the more common case where models achieve higher accuracy on textual variants compared to visual variants.
@@ 4 Cross-modality Circuit Analysis
Figure 3: Patching effects for Qwen2-7B-VL for the textual (left) and visual (right) counting task. We sum the patching effect across all model components for a specific position and layer. This reveals different patterns of component importance by position, motivating the separation of each circuit to three sub-circuits—data, query and generation.
To investigate the performance gap, we aim to pinpoint the differences between the circuits used for analogous visual and textual tasks. Inspired by Marr’s 3 levels of analysis (Marr & Poggio, 1976), we investigate both the structural and functional intersections between these circuits. We start by identifying the circuits responsible for each combination of model, task, and modality.
@@ 4.1 Vision and Language Circuits are Implemented by Different Components
Similarly to Kaduri et al. (2024), and motivated by Figure 3, we separate our prompts into three parts: data positions, query positions, and the answer generation position—the last token of the prompt. The positions are consistent within each modality, as detailed in Section 3, but can differ between modalities. Since circuit components are position-specific, we have to map the data, query, and generation positions between the two modalities in order to evaluate their overlap. Query positions and generation positions are easily mapped, since they consist of the same textual sequence for both textual and visual inputs (e.g., ”How many “banana” are in the image/sequence”?). A special case of alignment is in the data positions, where a single text position cannot be mapped to a single image patch. Thus, we do not consider data positions when calculating intersections for the data processing components (i.e., we only compare the layer and attention head index / MLP neuron index).
Consequently, we split each circuit to three sub-circuits, split by positions: data ($\mathbf{c}^{D}$), query ($\mathbf{c}^{Q}$) and generation ($\mathbf{c}^{G}$) sub-circuits, which each include the circuit components in the corresponding positions, such that $\mathbf{c}=\mathbf{c}^{D}\cup\mathbf{c}^{Q}\cup\mathbf{c}^{G}$.
@@
$\operatorname{IoU}\left(\mathbf{c}_{L},\mathbf{c}_{V}\right)=\frac{1}{2}\left(\frac{\left|\mathbf{c}_{L\to V}\cap\mathbf{c}_{V}\right|}{\left|\mathbf{c}_{L\to V}\cup\mathbf{c}_{V}\right|}+\frac{\left|\mathbf{c}_{L}\cap\mathbf{c}_{V\to L}\right|}{\left|\mathbf{c}_{L}\cup\mathbf{c}_{V\to L}\right|}\right)$ (3)
where $\mathbf{c}_{L}$, $\mathbf{c}_{V}$ are the textual and visual circuits, respectively. $\mathbf{c}_{L\to V}$ and $\mathbf{c}_{V\to L}$ are the textual and visual circuits with their positions mapped to the other modality, respectively. Due to the significant numerical imbalance between MLP neurons and attention heads, we calculate the intersection in eq. 3 separately for MLP neurons and attention heads, and average the results.
@@
$\operatorname{NIoU}\left(\mathbf{c}_{L},\mathbf{c}_{V}\right)=\frac{\operatorname{IoU}\left(\mathbf{c}_{L},\mathbf{c}_{V}\right)-\operatorname{IoU}\left(\mathbf{c}_{LV},\mathbf{c}_{RV}\right)}{1.0-\operatorname{IoU}\left(\mathbf{c}_{LV},\mathbf{c}_{RV}\right)}$ (4)
@@
The normalized IoU results (Figure 5) reveal varying degrees of intersection across sub-circuits. In data positions, the modality-specific sub-circuits exhibit virtually no intersection across tasks and models—an expected finding given their specialized function in processing distinct data token distributions. In query positions, there is low overlap between modalities (12% on average), despite operating on identical input query tokens. The generation position sub-circuits show moderately higher intersection rates (38% on average), reflecting these components’ shared role in promoting consistent answers across modalities. Overall, these results indicate that while some components are important in both visual and textual tasks, the different modalities utilize relatively disjoint circuits.
@@
Despite being largely structurally disjoint, two distinct component sets may still implement similar functionality (Hanna et al., 2025). To measure the shared functionality of the textual and visual circuits, we build upon the partitioning of each circuit to three sub-circuits (Section 4.1). We investigate whether these sub-circuits are functionally interchangeable between the textual and visual circuits—for instance, can we replace $\mathbf{c}^{Q}_{L}$, the sub-circuit in the query positions of the language circuit, with $\mathbf{c}^{Q}_{V}$, the sub-circuit in the query positions of the visual circuit, while maintaining high circuit faithfulness? Such interchangeability would indicate shared functionality. Formally, we interchange the query sub-circuit between modalities and measure the average faithfulness (as in eq. 2):
$\frac{1}{2}\left(\operatorname{F}(\mathbf{c}_{L}^{D}\cup\mathbf{c}_{V\to L}^{Q}\cup\mathbf{c}_{L}^{G},\mathrm{T_{L}})+\operatorname{F}(\mathbf{c}_{V}^{D}\cup\mathbf{c}_{L\to V}^{Q}\cup\mathbf{c}_{V}^{G},\mathrm{T_{V}})\right)$ (5)
We apply this process to data and generation sub-circuits as well. When interchanging data sub-circuits, we ignore positional information (as in Section 4.1) and treat components as active across all data positions. As in eq. 4, we normalize these scores using a random baseline as a lower bound and the original faithfulness as an upper bound. The random baseline is measured by interchanging a sub-circuit with an identically-sized sub-circuit with random components, and the upper bound is the original faithfulness of the entire circuit ($\mathbf{c}_{L}$ or $\mathbf{c}_{V}$).
@@
Our results (Figure 6) indicate that in the query and generation positions, textual and visual sub-circuits are interchangeable. In query positions, the visual sub-circuit can be replaced with a position-aligned textual sub-circuit (and vice versa), while maintaining high faithfulness to the entire model. This high interchange extends to the generation sub-circuits, showing a functional equivalence. Given the low structural overlap (12% and 38% on average in the query and generation positions, respectively, shown in Section 4.1), this implies that different components at these positions implement similar functionality, indicating a form of redundancy. A different pattern emerges when interchanging the sub-circuits at the data positions: across all tasks, the faithfulness drops to match the random baseline, indicating that the data sub-circuits $\mathbf{c}_{L}^{D}$ and $\mathbf{c}_{V}^{D}$ implement completely different functionality in textual and visual prompts. This indicates that the functional difference lies mainly in the components responsible for processing the data tokens, prior to being processed by the rest of the model.

### v3

@@ 1 Introduction
Recent years have seen major progress in building vision-and-language models (VLMs) that can handle both image and text inputs (Li et al., 2025). However, a significant performance gap persists between textual tasks and their visual counterparts (Fu et al., 2024; Wang et al., 2024a; Cohen et al., 2024). For example, models are considerably less accurate at counting objects in an image than at counting words in a list (Figure˜1a, Appendix˜C), or at identifying a winner in a board game when it is supplied as an image rather than as a textual description (Fu et al., 2024). Curiously, this performance gap persists even in VLMs with a decoder that is jointly trained from scratch on both modalities (e.g. Team et al. (2024)). To address the performance gap between analogous image and text tasks in VLMs, one should first understand its source.
@@
In this paper, we aim to explain the performance gap in VLMs between analogous visual and textual tasks from structural and functional perspectives. First, we create a dataset of five diverse tasks, each comprising pairs of analogous textual and visual prompt variants (Figure˜2). Each prompt consists of data, query, and generation positions, that contain the subject of the prompt (an image or a short text), the task description, and the last token position, respectively. For example, the data for an object counting task can be a text listing objects (“car flower banana…”) or an image with these objects, and the query for this task will ask “How many bananas are there?”. We use this dataset to extract and examine the circuits—the computational sub-graphs—responsible for performing each of the variants. We measure the degree of structural overlap between these modality-specific circuits, and find that vision and language tasks are performed by relatively disjoint VLM components (Figure˜1a), with an average of only 18% components shared between modalities.
@@
To analyze the functional differences of intermediate sub-circuits, we swap them between the two modalities, and measure how interchangeable they are for task performance (Figure˜1b). This reveals two main findings: first, despite being structurally disjoint, the sub-circuit that processes the query tokens, as well as the sub-circuit that generates the answer, are functionally equivalent between modalities—leading to similar performance when swapped. Second, there is a major functional difference in the processing of data tokens between modalities, resulting in much lower performance after swapping the relevant sub-circuit. Hence, we conclude that the difference in data processing is the leading cause behind the accuracy gap.
@@
We utilize these findings to bridge the performance gap, without any additional training, by automatically intervening in the model’s computation at test time. Specifically, we patch the visual token embeddings from later layers of the model—where these embeddings are more aligned to their textual counterparts—back to earlier layers (Figure˜1c). This results in visual processing that is more aligned with analogous textual prompts, leading to higher accuracy on visual prompts. Our results suggest that understanding the differences in the mechanisms between modalities is attainable, and provides a path to improving visual processing in VLMs.
@@ 2 Preliminaries
In this section we present our main unit of analysis—model circuits. We discuss how circuits for specific tasks are discovered and evaluated. These discovery and evaluation techniques are used in Section˜4 to identify the common and distinct sub-circuits for analogous tasks in different modalities.
@@ 2.1 Model Circuits
A circuit $\mathbf{c}$ is a minimal subset of interconnected model components that perform the computations required for a specific task or prompt (Elhage et al., 2021). In our analysis, we consider a circuit component to be either an entire attention head or an individual multilayer perceptron (MLP) neuron, at a specific output position. We follow the existing definition of MLP neurons (Geva et al., 2021; Nikankin et al., 2024), where the $n^{\text{th}}$ neuron in an MLP block is the combination of the $n^{\text{th}}$ row vector of the first MLP layer and the $n^{\text{th}}$ column vector of the second MLP layer. A full definition for MLP neurons and circuits is given in Appendix˜A.
@@ 2.2 Circuit Discovery
To identify circuits for both textual and visual task variants, we employ causal analysis techniques that score the importance of each component’s activation (Vig et al., 2020). For each textual or visual task, we divide its prompts into two subsets: one for discovery and another for evaluation (See Section˜B.3). From the discovery subset, for each prompt $p$ with answer $r$, we select a random counterfactual prompt $p^{\prime}$ that yields a different answer $r^{\prime}$. Ideally, we would find the importance of a component $u$ for a prompt $p$ by intervening on its activation $a_{u,p}$ with the counterfactual activation $a_{u,p^{\prime}}$, and calculate the difference in logits of $r$ and $r^{\prime}$ post-intervention (Wang et al., 2022), marked $\operatorname{LD}(r,r^{\prime}|\operatorname{do}(a_{u,p}=a_{u,p^{\prime}}))$. However, this process becomes unfeasible due to the large amount of model components. Instead, we approximate this effect by applying attribution patching (Nanda, 2022) with integrated gradients (AP-IG; Sundararajan et al., 2017; Hanna et al., 2024):
$\operatorname{E}(u,p,p^{\prime})=\left|\left(a_{u,p^{\prime}}-a_{u,p}\right)\frac{1}{k}\sum_{i=1}^{k}\frac{\partial\operatorname{LD}\left(r,r^{\prime}|\operatorname{do}\left(e=e^{\prime}+\frac{i}{k}\left(e-e^{\prime}\right)\right)\right)}{\partial a_{u,p}}\right|\ ,$ (1)
where $e$ and $e^{\prime}$ are the input embeddings for the prompts $p$ and $p^{\prime}$, respectively, and $k=5$ is the number of integration steps, following Hanna et al. (2024). This value is a first-order Taylor-series approximation to the logit difference, and can be calculated once, for all model components in parallel. The value is averaged across prompts and calculated per circuit component. A higher value indicates greater component importance for the analyzed task prompts. Given a score for each component, we construct a circuit $\mathbf{c}$ by selecting a specific percentage of the highest-scoring components.
@@ 2.3 Circuit Evaluation
We evaluate a circuit $\mathbf{c}$ by measuring its faithfulness (Wang et al., 2022)—the proportion of the full model’s task performance explained by the circuit. For each prompt $p$ with answer $r$ in the evaluation subset, we randomly choose a counterfactual prompt $p^{\prime}$ with answer $r^{\prime}$ and compute the activations of the model for $p^{\prime}$. We then pass $p$ through the model while ablating all non-circuit components using the pre-computed counterfactual activations, as per standard procedure (Mueller et al., 2025). The intervention’s effect is quantified by the logit difference between correct and counterfactual answers, denoted $\operatorname{LD}(r,r^{\prime}|\operatorname{do}(\forall u\notin\mathbf{c}:a_{u,p}=a_{u,p^{\prime}}))$. This effect is normalized to calculate the faithfulness of circuit $\mathbf{c}$ on task $\mathrm{T}$ (Mueller et al., 2025):
$\operatorname{F}(\mathbf{c},\mathrm{T})=\frac{1}{|\mathrm{T}|}\sum_{(p,p^{\prime},r,r^{\prime})\in\mathrm{T}}\frac{\operatorname{LD}(r,r^{\prime}|\operatorname{do}(\forall u\notin\mathbf{c}:a_{u,p}=a_{u,p^{\prime}}))-\operatorname{LD}_{\mathbf{M}}(r,r^{\prime})}{\operatorname{LD(r,r^{\prime})}-\operatorname{LD}_{\mathbf{M}}(r,r^{\prime})}\ ,$ (2)
@@ 3 Experimental Settings
To investigate the accuracy gap between textual and visual prompts in VLMs, we construct a dataset of five question-answering tasks, seen in Figure˜2. Each task consists of a query paired with data presented in one of two analogous formats: as either an image or a text. This analogous design ensures direct comparability—for every image-based prompt, we create an analogous text-based prompt that describes the same information. Our dataset contains the following tasks:
@@
To perform circuit discovery on these tasks, we enforce two constraints on the prompts. First, we only use prompts where the answer is a single word. Second, for consistency within a task, we align all prompts to a positional template, such that each position in each prompt always contains the same type of token in it. (For instance, in factual recall, the first four tokens contain the entity name.) Prompts that do not align well to the template are filtered out. These constraints are standard in circuit discovery (Vig et al., 2020; Mueller et al., 2025). Prompt examples per task and further details on prompt generation are provided in Section˜B.2.
@@
To ensure we identify meaningful circuits within these models, we follow standard procedures (Mueller et al., 2025) and verify high VLM performance (substantially above-chance accuracy) on each task. We report accuracies in Appendix˜C and focus on the more common case where models achieve higher accuracy on textual variants compared to visual variants.
@@ 4 Cross-modality Circuit Analysis
Figure 3: Patching effects for Qwen2-7B-VL for the textual (left) and visual (right) counting task. We sum the patching effect (described in Section˜2.2) across all model components for a specific position and layer. This reveals different patterns of component importance by position, motivating the separation of each circuit to three sub-circuits—data, query and generation.
To investigate the performance gap, we aim to pinpoint the differences between the circuits used for analogous visual and textual tasks. Namely, we analyze both the structural and functional intersection between these circuits. We start by identifying the circuits responsible for each combination of model, task, and modality.
@@ 4.1 Vision and Language Circuits are Implemented by Different Components
Similarly to Kaduri et al. (2024), and motivated by Figure˜3, we separate our prompts into three parts: data positions, query positions, and the answer generation position—the last token of the prompt. The positions are consistent within each modality, as detailed in Section˜3, but can differ between modalities. Since circuit components are position-specific, we have to map the positions between the two modalities in order to evaluate their overlap. Query positions and generation positions are easily mapped, since they consist of the same textual sequence for both textual and visual inputs (e.g., “How many ‘banana’ are in the image/sequence”?). A special case of alignment is in the data positions, where a single text position cannot be mapped to a single image patch. Thus, we do not consider data positions when calculating intersections for the data processing components (i.e., we only compare the layer and attention head index / MLP neuron index). This is formally defined in Section˜A.3.
Consequently, we split each circuit to three sub-circuits, split by positions: data ($\mathbf{c}^{D}$), query ($\mathbf{c}^{Q}$) and generation ($\mathbf{c}^{G}$) sub-circuits, which each include the circuit components in the corresponding positions, such that $\mathbf{c}=\mathbf{c}^{D}\cup\mathbf{c}^{Q}\cup\mathbf{c}^{G}$. A formal definition is given in Section˜A.2.
@@
$\operatorname{IoU}\left(\mathbf{c}_{L},\mathbf{c}_{V}\right)=\frac{1}{2}\left(\frac{\left|\mathbf{c}_{L\to V}\cap\mathbf{c}_{V}\right|}{\left|\mathbf{c}_{L\to V}\cup\mathbf{c}_{V}\right|}+\frac{\left|\mathbf{c}_{L}\cap\mathbf{c}_{V\to L}\right|}{\left|\mathbf{c}_{L}\cup\mathbf{c}_{V\to L}\right|}\right)\ ,$ (3)
where $\mathbf{c}_{L}$, $\mathbf{c}_{V}$ are the textual and visual circuits, respectively. $\mathbf{c}_{L\to V}$ and $\mathbf{c}_{V\to L}$ are the textual and visual circuits with their positions mapped to the other modality, respectively (formally defined in Section˜A.3). Due to the significant numerical imbalance between MLP neurons and attention heads, we calculate the intersection in eq.˜3 separately for MLP neurons and attention heads, and average the results.
@@
$\operatorname{NIoU}\left(\mathbf{c}_{L},\mathbf{c}_{V}\right)=\frac{\operatorname{IoU}\left(\mathbf{c}_{L},\mathbf{c}_{V}\right)-\operatorname{IoU}\left(\mathbf{c}_{LV},\mathbf{c}_{RV}\right)}{1.0-\operatorname{IoU}\left(\mathbf{c}_{LV},\mathbf{c}_{RV}\right)}\ .$ (4)
@@
The normalized IoU results (Figure˜5) reveal varying degrees of intersection across sub-circuits. In data positions, the modality-specific sub-circuits exhibit virtually no intersection across tasks and models—an expected finding given their specialized function in processing distinct data token distributions. In query positions, there is low overlap between modalities (12% on average), despite operating on identical input query tokens. The generation position sub-circuits show moderately higher intersection rates (38% on average), reflecting these components’ shared role in promoting consistent answers across modalities. Overall, these results indicate that while some components are important in both visual and textual tasks, the different modalities utilize relatively disjoint circuits.
@@
Despite being largely structurally disjoint, two distinct component sets may still implement similar functionality (Hanna et al., 2025). To measure the shared functionality of the textual and visual circuits, we build upon the partitioning of each circuit to three sub-circuits (Section˜4.1). We investigate whether these sub-circuits are functionally interchangeable between the textual and visual circuits—for instance, can we replace $\mathbf{c}^{Q}_{L}$, the sub-circuit in the query positions of the language circuit, with $\mathbf{c}^{Q}_{V}$, the sub-circuit in the query positions of the visual circuit, while maintaining high circuit faithfulness? Such interchangeability would indicate shared functionality. Formally, we interchange the query sub-circuit between modalities and measure the average faithfulness (as in eq.˜2):
$\frac{1}{2}\left(\operatorname{F}(\mathbf{c}_{L}^{D}\cup\mathbf{c}_{V\to L}^{Q}\cup\mathbf{c}_{L}^{G},\mathrm{T_{L}})+\operatorname{F}(\mathbf{c}_{V}^{D}\cup\mathbf{c}_{L\to V}^{Q}\cup\mathbf{c}_{V}^{G},\mathrm{T_{V}})\right)\ .$ (5)
We apply this process to data and generation sub-circuits as well. When interchanging data sub-circuits, we ignore positional information (as in Section˜4.1) and treat components as active across all data positions. As in eq.˜4, we normalize these scores using a random baseline as a lower bound and the original faithfulness as an upper bound. The random baseline is measured by interchanging a sub-circuit with an identically-sized sub-circuit with random components, and the upper bound is the original faithfulness of the entire circuit ($\mathbf{c}_{L}$ or $\mathbf{c}_{V}$).
@@
Our results (Figure˜6) indicate that in the query and generation positions, textual and visual sub-circuits are interchangeable. In query positions, the visual sub-circuit can be replaced with a position-aligned textual sub-circuit (and vice versa), while maintaining high faithfulness to the entire model. This high interchange extends to the generation sub-circuits, showing a functional equivalence. Given the low structural overlap (12% and 38% on average in the query and generation positions, respectively, shown in Section˜4.1), this implies that different components at these positions implement similar functionality, indicating a form of redundancy. A different pattern emerges when interchanging the sub-circuits at the data positions: across all tasks, the faithfulness drops to match the random baseline, indicating that the data sub-circuits $\mathbf{c}_{L}^{D}$ and $\mathbf{c}_{V}^{D}$ implement completely different functionality in textual and visual prompts. This indicates that the functional difference lies mainly in the components responsible for processing the data tokens, prior to being processed by the rest of the model.